OpenStudy (anonymous):

Backhaus Beer Brewers (BBB) just announced that the current fiscal year's income statement reports its net income to be $1.2 million. BBB's marginal tax rate is 40 percent, and its interest expense for the year was $1.5 million. The company has $8.0 million of invested capital, of which 60 percent is debt. In addition, BBB tries to maintain a weighted average cost of capital (WACC) near 12 percent. A) Compute the operating income (EBIT) that BBB earned in the current year B) What is BBB's economic value added (EVA) for the current year?

OpenStudy (anonymous):

C) BBB has 500,000 shares of common stock outstanding. According to the EVA value you computed in question B, how much can BBB pay in dividends per share before the value of the firm would start to decrease? If BBB doesn't pay any dividends, what would you expect to happen to the value of the firm?

OpenStudy (anonymous):

A)Add back Interest and Taxes to Net Income. B)Use the formula for EVA. c)Use your logic on this one, it's pretty straightforward.

OpenStudy (ar43r):

is this calculation correct? A) ebit = net income + interest + taxes ebit = 1.2 + 1.5 + 0.8 = 3.5 r = 3.5/8 = 0.4375 B) eva = r*Capital - c*Capital where r = rate of return = ebit/capital invested c = wacc eva = 0.4375*8 - 0.12*8 = 2.54

OpenStudy (ar43r):

I'm not sure that my premises are correct. The calculation for C looks like this: C) supposing the company paid all earnings in dividends dividends paid per share = X 500.000*X - 2.54 = 0 X = 2540000/500000 = 5.08

OpenStudy (anonymous):

hey ar43r, the answer for a matched the answer indicated in the book, but the answer for b did not match. According to the book, the answer for b is supposed to be 1.14 million. I am not sure about c

OpenStudy (anonymous):

The answer for b is wrong because he forgot to take the NOPAT after taxes; he used the NOPAT. As for question c, his formula is right but the numbers are wrong.

OpenStudy (ar43r):

okay. have found another computation of eva. also for B) eva = nopat - wacc*capital where nopat = (1-tax rate) * ebit. (see http://en.wikipedia.org/wiki/NOPAT) napat = (1-0.4) * 3.5 = 2.1 eva = 2.1-0.12*8 = 1.14

OpenStudy (anonymous):

That is correct but you can still use your old formula; for the calculation of "r" use EBIT(1-T) and you will get the correct answer.

OpenStudy (ar43r):

thanx BruLee! did you also know the answer for C)?

OpenStudy (anonymous):

For c) what you did is right but you have to replace the EVA value with the new one to get correct answer.

OpenStudy (ar43r):

thx!